OpenStudy (anonymous):

Solve the system of equations by the substitution method. Choose one answer. a. (7, 2) b. (-8, 2) c. (-7, 1) d. no solution

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

Sorry! here it is!

OpenStudy (radar):

isolate x in the x+4y=-11 equation x=-11-4y now substitute this for x in the other equation 9(-11-4y)+3y=-66 -99-36y+3y=-66 -33y=-66+99 -33y=33 y=-1 x-4=-11 x=-7

OpenStudy (radar):

Now we need to verify as we have another answer. -7+4(-1)=-11 -11=-11 that one is o.k. 9(-7)+3(-1)=-66 -63-3=-66 -66=-66 It is a good bingo

OpenStudy (radar):

Did you follow with understandingWynWynn?